how to calculate the packing fraction and binding energy of sodium in Me? if the mass of proton is 1.00782u ,mass of neutrons 1.00866u and mass of sodium is 22.989u

Grade:

2 Answers

Sunil Kumar FP
9 years ago
packing fraction of sodium will be given by the formula
=volume of atom in aunit cell/volume of the unit cell

binding energy of sodium is given by=
E = (delta)m c^2

(delta)m = 11*mp +12*mn+11*me-m(sodium 23)
= 11*1.007276 amu +12*1.008665 amu - 22.9897692809 amu
= 0.2002813191 * 1.66053892 × 10^-27 kilograms
= 3.3257493*10^-28 Kg

E = 3.3257493*10^-28 * 299 792 458 ^ 2
=2.9890344e-11j
=186.814647778 MeV
Sunil Kumar FP
9 years ago

packing fraction of sodium will be given by the formula
=volume of atom in aunit cell/volume of the unit cell

binding energy of sodium is given by= E = (delta)m c^2 (delta)m =
11*mp +12*mn+11*me-m(sodium 23) =
11*1.007276 amu - 22.9897692809 amu = 0.2002813191 * 1.66053892 × 10^-27 kilograms = 3.3257493*10^-28 Kg
E = 3.3257493*10^-28 * 299 792 458 ^ 2 =2.9890344e-11j =
186.814647778 MeV